how do you write the equation in slope-intercept form for line through the two points. (1,-2) and (3,10)
Found 2 solutions by rfer, amarjeeth123:
Answer by rfer(16322)
Answer by amarjeeth123(569)

Slope of the line m=(y2-y1)/(x2-x1)=(10+2)/(3-1)=12/2=6
Let the equation of the line be y=mx+c where m is the slope
Substituting the slope we get y=6x+c
Since it passes through (3,10) we get 10=18+c
c=10-18
c=-8
The equation of the line is y=6x-8